OpenStudy (anonymous):

imagine that the government of a small community decides to give a total of w, distributed equally, to all its citizens. Suppose that each month each citizen saves a fraction p of his or her new wealth and spends the remaining 1-p in the community. Assume no money leaves or enters the community, and all the money is redistributed throughout the community. a. if this cycle is continued for many months, how much money is ultimately spent? specifically by what factor is the initial investment of w increased. b. evaluate the limits p-->0 and p-->1 and interpret their meanings.

OpenStudy (anonymous):

w+w(1-p)+w(1-p)^2… a/1-r w/(1-1-p)

OpenStudy (anonymous):

@asnaseer

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@sirm3d @chihiroasleaf

OpenStudy (anonymous):

what i did at the start was the start of the problem having trouble with the rest.

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Gotta go to sleep, may have a look at this tomorrow.

OpenStudy (anonymous):

don't bother i'll have done by then.

OpenStudy (anonymous):

thanks anyways.

OpenStudy (anonymous):

let's see...

OpenStudy (anonymous):

does it give an interest rate at which the money is saved?

OpenStudy (anonymous):

does the government give each citizen w each month? or only once?

OpenStudy (anonymous):

once, and interest rate.

OpenStudy (anonymous):

no interest rate.

OpenStudy (anonymous):

is provided.

OpenStudy (anonymous):

p.s. i'm ridiculously tired so please excuse my mistakes and otherwise terrible answers.

OpenStudy (anonymous):

so... the citizens get one amount of money, w

OpenStudy (anonymous):

yeps.

OpenStudy (anonymous):

and they spend a fraction p of it every time

OpenStudy (anonymous):

yeps, pretty much.

OpenStudy (anonymous):

so then the amount of money they have left would be\[$=(1-p)^t\] where t is time in months

OpenStudy (anonymous):

if it goes on forever, they spend all the money

OpenStudy (anonymous):

w is not increased

OpenStudy (anonymous):

so the question doesn't really make sense

OpenStudy (anonymous):

unless p is such that they don't spend any money

OpenStudy (anonymous):

you forgot something... w(n+1)=wn(1-p)^t

OpenStudy (anonymous):

() denote subscripts

OpenStudy (anonymous):

that's about what I said with the equation $= just more accurate. unless the government is giving more money out after the start

OpenStudy (anonymous):